HanAkoSolo (jamierox4ev3r):

If sin(x) = 1/3 and sec(y) = 5/4, where x and y lie between 0 and π/2, evaluate sin(x + y).

HanAkoSolo (jamierox4ev3r):

@ganeshie8 I actually have no idea what's going on here, so if you're online, I would love your help. thanks

ganeshie8 (ganeshie8):

recall the angle sum identity for sin : \[\sin(x+y)=\sin x\cos y + \cos x\sin y\]

ganeshie8 (ganeshie8):

basically you need to find those four pieces : \(\sin x, \cos x, ~~\sin y, \cos y\) then plug them in

HanAkoSolo (jamierox4ev3r):

Alright, so I'm assuming that it would look a little like this: \(\sin (x + y) = (\frac{1}{3} )(\frac{5}{4})+ \sin(y)\cos(x) \)

HanAkoSolo (jamierox4ev3r):

that's all we appear to know at this moment in time

ganeshie8 (ganeshie8):

you have the correct idea but we don't really know the value of \(\cos y\) yet, what we're given is \(\sec y\) and not \(\cos y\)

ganeshie8 (ganeshie8):

so we need to find \(\cos y\), \(\sin y\) and \(\cos x\) somehow from the given info

HanAkoSolo (jamierox4ev3r):

oh whoops, completely forgot about the secant ;-;

HanAkoSolo (jamierox4ev3r):

riiight

ganeshie8 (ganeshie8):

its easy, lets find them

HanAkoSolo (jamierox4ev3r):

Fair enough. Let's do it

ganeshie8 (ganeshie8):

do you like triangles we're gonna need bunch of them

HanAkoSolo (jamierox4ev3r):

triangles are good with me

ganeshie8 (ganeshie8):